### 15 Dec 2023 v3

# Fine-Grained Image-Text Alignment in Medical Imaging Enables Cyclic Image-Report Generation

Abstract
Fine-grained vision-language models (VLM) have been widely used for inter-modality local alignment between the predefined fixed patches and textual words. However, in the medical analysis, lesions exhibit varying sizes and positions, and the fixed patches may cause incomplete representations of lesions. Moreover, these methods provide explainability by using heatmaps to show the general potential image areas associated with texts rather than specific regions, making their explanations not explicit and specific enough. To address these issues, we propose a novel Adaptive patch-word Matching (AdaMatch) model to correlate chest X-ray (CXR) image regions with words in medical reports and apply it to CXR-report generation to provide explainability for the generation process. AdaMatch exploits the fine-grained relation between adaptive patches and words to provide explanations of specific image regions with corresponding words. To capture the abnormal regions of varying sizes and positions, we introduce the Adaptive Patch extraction (AdaPatch) module to acquire the adaptive patches for these regions adaptively. In order to provide explicit explainability for CXR-report generation task, we propose an AdaMatch-based bidirectional large language model for Cyclic CXR-report generation (AdaMatch-Cyclic). It employs the AdaMatch to obtain the keywords for CXR images and ‘keypatches’ for medical reports as hints to guide CXR-report generation. Extensive experiments on two publicly available CXR datasets prove the effectiveness of our method and its superior performance to existing methods.
1 Introduction
Inter-modality alignment, such as vision and language, has been an important task with growing interests in the field of computer vision, especially with the recent advancement in representation learning [34]. Technologies like contrastive learning and self-supervised learning have dramatically improved state-of-the-art alignment performance. Recent vision-language models (VLMs) demonstrate two approaches: global contrastive alignment, which integrates images and texts at a global level [34, 19, 17, 44, 46], and local alignment, focusing on detailed connections between visual objects and textual words [6, 25, 24, 52, 21, 50], as illustrated in Fig. 1.
Figure 1: Current vision-language models (VLM) achieve (a) global alignment and (b) local alignment by matching overall visual with textual features, and aligning patch with word features, respectively. (c) To exploit the relation between textual words and abnormal patches with different sizes, the proposed AdaMatch obtains adaptive patch features and aligns them with word features.
Current VLMs with local alignment either adopt the pre-trained object detector to extract region-of-interest (ROI) features from images and match the corresponding object features with textual words [6, 25, 24, 52]; or align the visual token from each patch and the textual token into the same embedding space [21, 50, 18, 41]. The former highly relies on the quality of the object detector and its predefined classes, which is less generalizable to new domains. The latter family of methods learns the alignment in a more automatic and data-driven manner. However, most of these methods depend on pre-defined patch size and position (e.g., grids) across images. In most challenging cases, such as the analysis of medical image analysis, lesions can exhibit a wide range of shapes, sizes, and positions. A fixed definition of image patches can lead to incomplete or ambiguous representations of the key imaging abnormalities. Hence, it is highly needed to adaptively exploit the fine-grained relationship between image embeddings derived from a more flexible patching scheme and textual embeddings.
Another challenge in the current local medical VLMs lies in their explainability: it is generally difficult to delineate the image-text relationship learned by the model, especially for the current medical VLMs. Current solutions to provide such explanations in medical VLMs leverage the attention maps from the intermediate layer to visualize the location of the abnormalities [32, 16, 45]. Other methods  [40, 5, 28] utilize network gradients such as Grad-CAM [38] to generate the heatmaps according to lesion types based on ground-truth reports. However, both maps can only show the general potential areas associated with the corresponding text data rather than clearly pinpointing a specific region. In addition, gradient-based methods need ground-truth reports, prohibiting them from functioning correctly beyond training data. It is thus highly necessary to develop a mechanism that could provide explicit and specific explanations of the input image or text during inference time.
Figure 2: The overview of the proposed methods. (a) Adaptive patch-word Matching (AdaMatch) model. (b) AdaMatch-based bidirectional large language model (LLM) for cyclic CXR-report generation (AdaMatch-Cyclic).
To address these two challenges above, we propose a novel Adaptive patch-word Matching (AdaMatch) model to match fine-grained image regions of various sizes and positions with textual data. AdaMatch introduces an image encoder with multiple Adaptive Patch extraction (AdaPatch) modules to adaptively acquire the patches associated with certain text tokens. It then performs patch-word alignment based on contrastive learning. AdaMatch is specifically developed in the context of aligning radiology images (chest X-ray, CXR) and their corresponding radiology reports with the capability of achieving cyclic (CXR-to-report and report-to-CXR) generation based on the learned alignment. Our premise is that such a cyclic generation task would serve as the best use case and evaluation criterion for the desired fine-grained alignment. Also, the fine-grained cyclic generation between CXR and report will provide a natural explainability for how the model aligns the two modalities: for any given text token, we can visualize its matching imaging manifestation; and for any image region within a CXR image, we can tell the type of lesion or anatomical region it belongs to.
To implement the cyclic CXR-report generation, we propose an AdaMatch-based bidirectional model (AdaMatch-Cyclic). AdaMatch-Cyclic employs AdaMatch to identify the keywords for CXR images and ‘keypatches’ for medical reports to guide the generation tasks. Since the potential keywords for CXR images cover a wide range and ground-truth reports cannot be used during inference, we predefine a textual codebook with the most common entities from medical reports as prior knowledge during fine-grained alignment. With the textual codebook, AdaMatch aligns it with the adaptive patches to obtain matched keywords as guidance for report generation. Next, a VQ-GAN model encodes the CXR image into image tokens, and a Large Language Model (LLM) takes image tokens, the matched keywords, and the instructions as input to generate the medical reports. Similarly, we also build a visual codebook with the most commonly seen patches as ‘keypatches’, and use AdaMatch to obtain the matched keypatches from given text reports as hints for CXR generation. With medical reports, matched keypatches, and instructions, LLM outputs the image tokens, and then the VQ-GAN model decodes them into the generated CXR image. Our contributions are summarized as follows:
  • To exploit the fine-grained relation between the CXR image patches with words of medical reports, we propose an Adaptive patch-word Matching (AdaMatch) model to obtain adaptive patches for abnormal regions and perform alignment between them and texts in medical reports.
  • We devise an AdaMatch-based bidirectional LLM for Cyclic CXR-report generation (AdaMatch-Cyclic) to facilitate the generation between CXR and reports. Moreover, we build the textual and visual codebook to utilize AdaMatch to extract useful keywords and keypatches for the report and CXR generation, respectively.
  • Experiments on two publicly available chest X-ray datasets demonstrate the effectiveness of our method and its superior performance over state-of-art methods.
2 Related Works
2.1 Fine-Grained Vision-Language Models
Recently, several fine-grained vision-language models (VLM) [6, 25, 24, 52, 21, 50, 18, 41] achieve the local alignment by exploiting the fine-grained relation between visual objects and textual words. Some methods [6, 25, 24, 52] employ the pre-trained object detector to obtain region-of-interest (ROI) features from images and align the object features with textual features, and others [21, 50, 41, 18] aim to locally align the fixed patches with the textual words. For instance, Li et al. [24] utilizes Faster-RCNN [36] to detect visual objects and match ROI embeddings with textual embeddings. Moreover, FILIP [50] achieves cross-modal interaction by maximizing the similarity between the visual tokens of fixed patches and textual tokens. The former requires a precise object detector, and the latter primarily learns the relation between fixed patches inside the predefined grid and words. Nevertheless, when lung lesions differ in size, shape, and position, these methods may split them into separate patches, leading to incomplete semantic information of patches. Thus, we devise an Adaptive patch-word Matching (AdaMatch) model to adaptively exploit the fine-grained relation between flexible patches and textual words.
2.2 CXR-Report Generation
Medical VLMs are widely used in downstream tasks for chest radiographs, including CXR-to-report generation [7, 8, 48, 42, 39, 47] and report-to-CXR generation [37, 3, 2, 22, 23] tasks. CXR-to-report generation has been investigated to reduce the burden of radiologists and promote clinical automation. For instance, Chen et al. [8] introduces a cross-modal memory network with shared memory to align images with texts to promote report generation performance. In report-to-CXR generation task, previous methods generate high-quality and annotated CXR images from medical reports, to provide more training data and avoid privacy laws and considerations. Some diffusion-based methods [2, 22] take the impression section of reports as input, ignoring the finding section that includes detailed descriptions. To consider more details in reports, several transformer-based methods [22, 23] take both finding and impression sections as input to synthesize CXR. In this paper, we use the cyclic generation (i.e. CXR-to-report and report-to-CXR generation) as use case and evaluation criteria for our fine-grained alignment method (AdaMatch). We also design an AdaMatch-Cyclic to employ AdaMatch to improve the explainability of cyclic generation.
3 Methods
In Fig. 2, we propose an adaptive patch-word matching (AdaMatch) model to associate the CXR image regions with words in medical reports, and apply it to CXR-report generation to provide explainability. Given an input CXR image $x^{I}$, the image encoder with several Adaptive Patch extraction modules (AdaPatch) predicts the location and scale of multiple adaptive patches and processes the adaptive patch embeddings $z^{I}$. With $z^{I}$ and text embeddings $z^{T}$ extracted by text encoder for the medical report $x^{T}$, we compute the similarities $s^{I}$, $s^{T}$ between adaptive patches and text tokens to calculate the contrastive loss $L_{I2T},L_{T2I}$ to optimize AdaMatch model. For CXR-to-report generation, we utilize the frozen AdaMatch model to match a predefined textual codebook with the input CXR image $x^{I}$ to obtain matched keywords $k^{T}$ and feed a LLM with $k^{T}$, the instruction, and image tokens encoded by VQ-GAN from $x^{I}$ to generate a medical report $x^{T^{\prime}}$. Then, the AdaMatch model is used to match a predefined visual codebook with the generated report $x^{T^{\prime}}$ to acquire the matched keypatches $k^{I}$. With instruction, $x^{T^{\prime}}$ and $k^{I}$, LLM outputs image tokens of the generated CXR image $x^{I^{\prime}}$, which is optimized by $L_{text}$ and $L_{img}$.
3.1 Adaptive Patch-Word Matching (AdaMatch)
Current VLMs [6, 25, 24, 52, 21, 50] achieve fine-grained alignment between visual objects and textual words, but they may split lung lesions into separate fixed patches due to various shapes, sizes, and positions of lung lesions. Thus, we propose an Adaptive patch-word Matching (AdaMatch) to automatically locate the important regions, extract adaptive patches for these regions, and align them with corresponding textual words. As shown in Fig. 2 (a), AdaMatch includes an image encoder to extract adaptive patch features and a text encoder to encode textual features.
Image Encoder. To obtain the adaptive patches, the image encoder consists of four stages with feature maps of decreasing scales. Specifically, in the first stage, we first adopt the patch embedding module to split the CXR image $x^{I}\in\mathbb{R}^{H\times W\times C}$ into $N$ patches with fixed size $s\times s$ and project them to patch embeddings $z^{(i)}(1\leq i\leq N)$ through a fully connected layer $g$, $z^{(i)}=g([a^{(i,1)};\dots;a^{(i,s\times s)}])$, where $a^{(i,j)}$ indicates the image features for the pixel located at $q^{(i,j)}$ and []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation. Afterward, we pass the patch embeddings $z^{(i)}$ with the position embeddings $e_{pos}$ into a transformer encoder [43] $R$ to obtain the outputs, $z^{(i)}=R(z^{(i)}+e_{pos})$, where $R$ includes an attention layer and a feed-forward layer. In order to locate the potential lung lesions, we devise an Adaptive Patch Extraction module (AdaPatch) for the rest of stages. In AdaPatch, a fully connected layer is fed with patch embeddings $z^{(i)}$, and then the outputs are passed through four separate fully connected layers $f_{1},f_{2},f_{3},f_{4}$ to predict the offset $(\delta_{x}^{(i)},\delta_{y}^{(i)})$ and patch size $(s_{w}^{(i)},s_{h}^{(i)})$ for the adaptive patches, where the offset indicates the shift to the center of original fixed patches $(c^{(i)}_{x},c^{(i)}_{y})$,
$\delta_{x}^{(i)}=\small{Tanh}(f_{1}(z^{(i)})),\delta_{y}^{(i)}=\small{Tanh}(f_{2}(z^{(i)})),$ (1)
$s_{w}^{(i)}=\small{ReLU}(\small{Tanh}(f_{3}(z^{(i)})),s_{h}^{(i)}=\small{ReLU}(\small{Tanh}(f_{4}(z^{(i)})).$ (2)
With the offset and patch size, we compute the position of left-top $(a^{(i)}_{x},a^{(i)}_{y})$ and right-bottom corners $(b^{(i)}_{x},b^{(i)}_{y})$ for each adaptive patch,
$a^{(i)}_{x}=c^{(i)}_{x}+\delta_{x}^{(i)}-\frac{s_{w}^{(i)}}{2},a^{(i)}_{y}=c^{(i)}_{y}+\delta_{y}^{(i)}-\frac{s_{h}^{(i)}}{2},$ (3)
$b^{(i)}_{x}=c^{(i)}_{x}+\delta_{x}^{(i)}+\frac{s_{w}^{(i)}}{2},b^{(i)}_{y}=c^{(i)}_{y}+\delta_{y}^{(i)}+\frac{s_{h}^{(i)}}{2},$ (4)
and uniformly sample $m\times m$ feature points inside the patches. Since the coordinates may be fractional, bilinear interpolation is used to obtain the sampled feature points $\left\{\hat{p}^{(j)}\right\}_{1\leq j\leq m\times m}$. The embeddings of all sampled points are flattened and fed into a fully connected layer $f_{5}$ to obtain patch embeddings, $z^{(i)}=f_{5}([\hat{p}^{(i,1)};\dots;\hat{p}^{(i,m\times m)}])$. Finally, we pass $z^{(i)}$ with position embeddings $e_{pos}$ into a transformer encoder $R$. The final adaptive patch embeddings $z^{I}\in\mathbb{R}^{N\times d}$ are the ensemble of $z^{(i)}(1\leq i\leq N)$, $z^{I}=E^{I}(x^{I})=\left\{z^{I}_{1},\cdots,z^{I}_{N}\right\},$ where $E^{I}$, $x^{I}$, and $d$ denote the image encoder, the input CXR image, and the dimension of adaptive patch embeddings, respectively.
Text Encoder. We adopt the BioClinicalBERT [1] as the text encoder $E^{T}$ to encode the medical report $x^{T}$ into text embeddings $z^{T}$. To make $z^{T}$ with the same dimension as $z^{I}$, we feed a fully connected layer with $z^{T}$ to reduce its dimension, $z^{T}=E^{T}(x^{T})=\left\{z^{T}_{1},\cdots,z^{T}_{K}\right\}(z^{T}\in\mathbb{R}^{K\times d})$, where $K$ denotes the number of text tokens.
Patch-Word Alignment. In order to exploit the relation between the adaptive patches and textual tokens, we perform fine-grained contrastive representation learning to achieve patch-word alignment. Concretely, for the $i$-th CXR image $x^{I}_{i}$ and $j$-th medical report $x^{T}_{j}$, we first compute the similarities between all the adaptive patch embeddings $z^{I}_{n}(1\leq n\leq N)$ and all the text embeddings $z^{T}_{k}(1\leq k\leq K)$, and use the largest similarity $\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ as the patch-word maximum similarity for $n$-th adaptive patch embedding. Then, the patch-word maximum similarities for all the adaptive patch embeddings are averaged as the similarity $s^{I}_{i,j}$ of the $i$-th CXR image to the $j$-th medical report,
$s^{I}_{i,j}=\frac{1}{N}\displaystyle\sum_{n=1}^{N}(z^{I}_{n})^{\top}z^{T}_{m_{n}^{I}},$ (5)
where $m_{n}^{I}=\arg\max_{(1\leq k\leq K)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. Similarly, the similarity of the $j$-th medical report to the $i$-th CXR image is defined as,
$s^{T}_{i,j}=\frac{1}{K}\displaystyle\sum_{k=1}^{K}(z^{I}_{m_{k}^{T}})^{\top}z^{T}_{k},$ (6)
where $m_{k}^{T}=\arg\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$. We exclude the padded textual tokens when computing the similarity.
With the cross-modal similarities $s^{I}_{i,j}$ and $s^{T}_{i,j}$ for the $i$-th CXR image and $j$-th medical report, we compute the adaptive patch-word contrastive loss $L_{i}^{I}$ for $i$-th CXR image $x_{i}^{I}$,
$L_{i}^{I2T}(x_{i}^{I},\left\{x_{j}^{T}\right\}_{j=1}^{b})=-\frac{1}{b}\log\frac{exp(s^{I}_{i,i}/\tau)}{\textstyle\sum_{j}{exp(s_{i,j}^{I}/\tau)}},$ (7)
where $b$ denotes the batch size, $\tau$ represents temperature hyperparameter and $\left\{x_{i}^{I},x_{i}^{T}\right\}$ indicates the positive CXR-report pair. Similarly, the adaptive word-patch contrastive loss $L_{i}^{T}$ for $i$-th medical report $x_{i}^{T}$ is formulated as,
$L_{i}^{T2I}(x_{i}^{T},\left\{x_{j}^{I}\right\}_{j=1}^{b})=-\frac{1}{b}\log\frac{exp(s^{T}_{i,i}/\tau)}{\textstyle\sum_{j}{exp(s_{j,i}^{T}/\tau)}}.$ (8)
The final contrastive loss for a mini-batch is calculated by,
$L=\frac{1}{2}\sum_{i=1}^{b}{(L_{i}^{I2T}+L_{i}^{T2I})}.$ (9)
With the final contrastive loss, the image and text encoders are optimized to automatically locate the important patches with varied sizes and exploit the fine-grained relation between the patches and words.
3.2 AdaMatch-based LLM for Cyclic CXR-Report Generation (AdaMatch-Cyclic)
To provide explicit explainability for CXR-report generation task, we propose an AdaMatch-based bidirectional LLM for the cyclic CXR-report generation (AdaMatch-Cyclic) by locating the potential lesions in CXR images and visualizing the appearance of description in reports to guide the generation process, as depicted in Fig. 2 (b).
Figure 3: Instruction data example for CXR-to-report generation.
3.2.1 CXR-to-Report Generation
In CXR-to-report generation, we employ the AdaMatch to match a predefined textual codebook with the CXR image to obtain the matched keywords as hints to guide LLM to generate reports.
Building textual codebook. Specifically, to establish a textual codebook, we use a pre-trained BioEN [35] to extract the related entities from the medical reports in the training set, where the related entities are divided into four entity groups, i.e., biological structure, detailed description, disease disorder, and sign symptom. Next, we compute the frequency of each entity and pick top $\kappa_{0}$ entities for each entity group as keywords in the textual codebook.
Keywords Extraction. With the textual codebook, we employ the frozen AdaMatch model to match the keywords from the textual codebook with the adaptive patches of CXR images, and obtain a patch-word similarity matrix $s^{P}\in\mathbb{R}^{M\times N}$ between the keyword tokens and adaptive patch embeddings, where $M$ and $N$ denote the number of keyword tokens and adaptive patches. To extract the most matched keywords for each adaptive patch, we rank the patch-word similarity along the dimension of keyword tokens, obtain the top $\kappa_{1}$ patch-word similarities for each adaptive patch, and extract the corresponding keywords $k^{T}$. The matched keywords can explain potential lesions in each adaptive patch, to assist LLM in generating medical reports.
Figure 4: Instruction data example for report-to-CXR generation.
Instruction tuning LLM. After keywords extraction, we use a frozen VQ-GAN [13] encoder $E$ to encode the input CXR image $x^{I}$ to the quantized image latent vectors as image tokens $E(x^{I})$. For CXR-to-report generation, the dolly-v2-3b [9] model is adopted as the bidirectional LLM, which is pre-trained on the instruction-following dataset. When fine-tuning LLM, we convert CXR-to-report generation dataset in instruction-following format. An example of instruction data for CXR-to-report generation is shown in Fig. 3, including the instruction, input, keywords, and response parts, where input is the image tokens of the input CXR image, keywords are extracted by AdaMatch, and response is the ground-truth medical report. During training, the response part is hidden and the LLM learns to generate it in autoregressive manner.
By adopting AdaMatch-Cyclic for CXR-to-report generation, we can locate the potential lesion in CXR images, to provide more explainability for the generation process.
3.2.2 Report-to-CXR Generation
In report-to-CXR generation, we leverage the AdaMatch to match a predefined visual codebook with the generated reports $x^{T^{\prime}}$ to acquire matched keypatches as guidance for LLM to synthesize CXR images, where ‘keypatches’ are important image patches related to generated reports.
Building visual codebook. Concretely, we first construct a visual codebook with the most common adaptive patches as keypatches. To collect the most common adaptive patches for CXR images in the training set, the AdaMatch is utilized to match the adaptive patches of CXR images with textual tokens of medical reports to obtain top $\kappa_{2}$ CXR-report pairs with the highest report-to-CXR similarities $s^{T}$. For each CXR-report pair, we compute the word-patch maximum similarity $\max_{(1\leq n\leq N)}{{(z^{I}_{n})}^{\top}z^{T}_{k}}$ for each textual token, rank the word-patch maximum similarities, and extract the adaptive patches for top $\kappa_{3}$ similarities as keypatches in the visual codebook. Each keypatch includes its adaptive patches and the corresponding features.
Keypatches Extraction. With the visual codebook, the frozen AdaMatch matches the features of keypatches in the visual codebook with textual tokens of the generated report to acquire the word-patch similarity matrix $s^{W}\in\mathbb{R}^{(\kappa_{2}\times\kappa_{3})\times K}$, where $K$ denotes the number of textual tokens for the generated report. To obtain keypatches related to the generated report, we rank the word-patch similarity along the dimension of keypatches, obtain the top $\kappa_{4}$ word-patch similarity for each textual token, and extract the features of corresponding keypatches $k^{I}$.
Instruction tuning LLM. After keypatches extraction, we employ a frozen VQ-GAN encoder to convert the matched keypatches $k^{I}$ into image tokens $E(k^{I})$, and feed the pre-trained LLM with the instruction, the generated report, and image tokens of keypatches in the instruction-following format, as shown in Fig. 4. Then, LLM predicts the image tokens and the VQ-GAN decoder decodes the image tokens into the corresponding generated CXR image $x^{I^{\prime}}$.
AdaMatch-Cyclic allows us to interpret generated reports with matched keypatches, thereby providing explainability to the generation procedure.
3.2.3 Overall Objective
To optimize AdaMatch-Cyclic, we apply the standard language modeling objective for both CXR-to-report and report-to-CXR generation. For CXR-to-report generation, LLM is fed with the instruction $I^{t}$, image tokens of input CXR image $E(x^{I})$, and matched keywords $k^{T}$ in instruction-following format, and LLM aims to generate the ground-truth medical report of response part autoregressively. We compute the $P(u_{k})$ conditional probability of the $k$-th token, $P(u_{k})=\mathcal{LLM}(I^{t},E(x^{I}),k^{T})$, where $k$ is the token index after the response key (### Response:). The report generation loss $L_{text}$ for the response area is calculated by,
$L_{text}=\sum_{i=k}^{n}-\log P(u_{i}|u_{1},u_{2},\cdots,u_{i-1}),$ (10)
where $[u_{1},u_{2},\cdots,u_{i-1}]$ denotes the tokenized texts before the response part, and $n$ represents the maximum length of output tokens.
Figure 5: Qualitative comparison with existing methods in CXR-to-report generation on the MIMIC-CXR (1st row) and OpenI datasets (2nd row). The texts in different colors show similar meanings. The keywords on the right are obtained from AdaMatch model.
Similarly, for report-to-CXR generation, we feed the LLM with the instruction $I^{v}$, the generated report $x^{T^{\prime}}$, and the image tokens of keypatches $E(k^{I})$ in the instruction-following format, and obtain the conditional probability of $k$-th token $w_{k}$, $P(w_{k})=\mathcal{LLM}(I^{v},x^{T^{\prime}},E(k^{I}))$. The CXR generation loss $L_{img}$ is defined as:
$L_{img}=\sum_{i=k}^{n}-\log P(w_{i}|w_{1},w_{2},\cdots,w_{i-1}),$ (11)
With $L_{text}$ and $L_{img}$, the LLM can implement the bidirectional CXR and report generation according to instructions.
Table 1: CXR-to-report generation performance in comparison to existing methods on MIMIC-CXR and OpenI datasets.
MIMIC-CXR OpenI
Methods BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
R2Gen [7] (2020) 0.3553 0.2232 0.1523 0.1038 0.1412 0.2784 0.3992 0.2407 0.1518 0.0973 0.1390 0.3052
R2GenCMN [8] (2021) 0.3719 0.2332 0.1538 0.1053 0.1501 0.2827 0.4091 0.2493 0.1594 0.1045 0.1509 0.3181
Yang et al. [48] (2021) 0.3585 0.2266 0.1550 0.1021 0.1425 0.2788 0.3833 0.2409 0.1598 0.1078 0.1457 0.3293
XProNet [42] (2022) 0.3532 0.2212 0.1498 0.1052 0.1415 0.2811 0.4114 0.2502 0.1598 0.1045 0.1457 0.3240
ITHN [39] (2023) 0.3623 0.2128 0.1402 0.0992 0.1488 0.2622 0.2661 0.1516 0.0976 0.0663 0.1561 0.2617
M2KT [47] (2023) 0.3661 0.2192 0.1465 0.1044 0.1528 0.2673 0.2559 0.1381 0.0819 0.0523 0.1468 0.2439
AdaMatch-Cyclic 0.3793 0.2346 0.1540 0.1060 0.1625 0.2859 0.4161 0.3002 0.2073 0.1446 0.1621 0.3656
4 Experiments
4.1 Experiment Setting
Datasets. We conduct our experiments on two main publicly available chest X-ray datasets, i.e. MIMIC-CXR [20] and OpenI [11] datasets. MIMIC-CXR dataset is the largest public dataset for chest radiographs, including 473,057 chest X-ray images and 206,563 reports from 63,478 patients. We follow the official splits, i.e. 368,960 instances for training, 2,991 for validation, and 5,159 for testing. OpenI is a chest X-ray dataset with 3,684 report-image pairs, including 2,912 pairs for training and 772 pairs for testing. Different from existing methods [7, 8, 48, 47], we adopt both finding and impression sections as medical reports and remove the reports with less than 3 tokens for two datasets. There is no overlap of patients among different subsets. The CXR images with various sizes are all resized to $256\times 256$ pixels.
Implementation Details We adopt the PVT-medium [43] pre-trained with the disease classification task as the image encoder with AdaPatch in stage 2, 3 and 4. The AdaMatch is optimized with LAMB optimizer [51] and cosine learning rate scheduler [29]. The learning rate is 6e-3. The maximum training epoch is 15 and the per GPU batch size is 112. In AdaMatch-Cyclic, we adopt the VQ-GAN [13] models pre-trained on the MIMIC-CXR and OpenI datasets, respectively, and the dolly-v2-3b [9] model as pre-trained LLM. LLM has 5,1845 token types with 5,0821 token types for text tokens and 1,024 token types for image tokens. LLM is optimized by the AdamW [30] optimizer with a learning rate of 5e-6 and per GPU batch size of 24. The training epoch is 5. 8 Nvidia A100 40GB GPUs are used for all experiments. For more details, see supplementary materials.
Evaluation Metrics We evaluate the CXR-to-report generation performance with BLEU [33], METEOR [12], and ROUGE-L [27]. The FID [15] and NIQE [31] are utilized to evaluate report-to-CXR generation. The performance of CXR-to-report and report-to-CXR retrieval is assessed through the exact report in the top K retrieved reports for a given CXR image (R@K, K=$\left\{1,5,10\right\}$).
4.2 Comparison with State-of-the-Arts.
4.2.1 CXR-to-Report Generation
We verify the effectiveness of AdaMatch-Cyclic on the MIMIC-CXR and OpenI datasets for report generation. Since previous methods [7, 8, 48, 42, 39, 47] are mainly trained on the medical reports with the finding section, we reimplement previous methods on datasets with both finding and impression sections. In Table 1, our AdaMatch-Cyclic substantially surpasses the M2KT [47] by a significant margin with the BLEU-1 of 0.0132 in the MIMIC-CXR dataset. Moreover, we evaluate the OpenI dataset with the model trained on the MIMIC-CXR dataset. In terms of the zero-shot report generation, our AdaMatch-Cyclic achieves the best performance with the BLEU-1, METEOR, ROUGE-L of 0.4161, 0.1621, and 0.3656. To further compare with GPT-4V[49], we randomly select 50 cases from the test set of the MIMIC-CXR dataset. In Table 2, our AdaMatch-Cyclic significantly outperforms the GPT-4V by 0.1479 and 0.0892 of BLEU-1 and BLEU-4, indicating our superiority to existing methods.
Table 2: CXR-to-report generation performance in comparison to GPT-4V on 50 selected cases of the MIMIC-CXR dataset.
Methods BLEU-1 BLEU-2 BLEU-3 BLEU-4
R2Gen[7](2020) 0.3513 0.2174 0.1447 0.1025
R2GenCMN[8](2021) 0.3587 0.2200 0.1436 0.0969
Yang et al. [48](2021) 0.3596 0.2218 0.1481 0.1026
XProNet[42](2022) 0.3356 0.2071 0.1374 0.0941
ITHN[39](2023) 0.3301 0.1839 0.1121 0.0723
M2KT[47](2023) 0.3626 0.2123 0.1391 0.0957
GPT-4V[49](2023) 0.2275 0.0878 0.0378 0.0166
AdaMatch-Cyclic 0.3754 0.2303 0.1520 0.1058
Table 3: Report-to-CXR generation performance in comparison to existing methods on MIMIC-CXR and OpenI datasets.
MIMIC-CXR OpenI
[…]
Stable diffusion [37] 9.2334 3.7894 8.2946 6.3496
Chambon et al. [3] 8.2758 3.8871 5.8557 4.6534
RoentGen [2] 9.5411 3.8834 6.5675 4.9085
UniXGen [22] 6.7212 3.7125 11.9890 4.6610
LLM-CXR [23] 2.1788 3.5969 1.6597 3.8206
AdaMatch-Cyclic 1.0916 3.3931 1.5938 3.3096
To qualitatively evaluate report generation performance, we show two cases of generated reports in Fig. 5. In comparison to M2KT [47] and R2GenCMN [8], our AdaMatch-Cyclic can obtain the related keywords (e.g. scarring and atelectasis) and correctly describes the ‘osseous abnormality’ and ‘clips in the right upper quadrant of the abdomen’ in the first case, and point out ‘mild granulomatous’ in the second case, suggesting that our method can generate medical reports with higher quality than previous methods.
4.2.2 Report-to-CXR Generation
To quantitatively assess the performance, we compare our method with the text-to-image generation method [37] and some report-to-CXR generation methods [3, 2, 22, 23]. As listed in Table 3, AdaMatch-Cyclic achieves the best performance with 1.0916 and 1.5938 of FID scores on the MIMIC-CXR and OpenI datasets, respectively, indicating the superiority of our method over previous methods. To compare the high-level feature distribution of CXR images generated by different methods, we randomly select 1,000 cases from the test set, and apply the t-SNE visualization to the real and synthetic CXR images on the MIMIC-CXR dataset. As depicted in Fig. 6, synthetic CXR images generated by existing methods and real ones differ significantly from each other, while CXR images synthesized by AdaMatch-Cyclic almost overlap with the real CXR images. It implies that AdaMatch-Cyclic is more effective in translating medical reports to realistic CXR images compared to existing methods.
Figure 6: The t-SNE visualization of the real and synthetic CXR images on the MIMIC-CXR dataset.
Figure 7: The generated chest X-ray images of the MIMIC-CXR (1st row) and OpenI (2nd row) datasets with highlighted regions.
In Fig. 7, we show the generated CXR images of the MIMIC-CXR and OpenI datasets. In the first example, our AdaMatch-Cyclic can synthesize the ‘left and right pleural effusions’ as highlighted in the CXR images, while other methods [3, 2, 37] cannot generate the CXR images with ‘pleural effusions’. It suggests that our method can generate more realistic CXR images than existing methods according to the input report.
4.2.3 CXR-Report Retrieval
To prove the superiority of AdaMatch, we evaluate the CXR-to-report and report-to-CXR retrieval performance on the MIMIC-CXR datasets in comparison with existing methods. As listed in Table 4, our AdaMatch achieves the best retrieval performance, surpassing the LIMITR [10] by a large margin with R@1 of 11.77% and 13.48% for CXR-to-report and report-to-CXR retrieval, respectively, implying that our method is more effective in extracting distinct semantic features for CXR images and medical reports and is able to align their features correctly.
Table 4: CXR-report retrieval performance (%) in comparison to existing methods on the MIMIC-CXR dataset.
CXR-to-Report Report-to-CXR
Methods R@1 R@5 R@10 R@1 R@5 R@10
Fang et al. [14] (2015) 18.60 43.10 56.10 18.13 43.20 55.97
Chauhan et al. [4] (2020) 5.37 19.43 30.73 5.40 20.23 30.23
ConVIRT [53] (2020) 30.10 53.90 63.80 29.20 54.70 64.40
GLoRIA [16] (2021) 30.30 57.50 66.50 24.00 51.80 62.80
JoImTeR-Net [18] (2021) 18.93 46.20 58.67 19.07 45.27 58.50
MGCA [41] (2022) 25.80 51.90 62.10 27.90 51.20 61.60
LIMITR [10] (2023) 39.70 63.20 71.70 37.70 62.10 71.30
Motor [26] (2023) 10.96 31.93 42.90 12.00 33.10 44.32
AdaMatch 51.47 86.19 94.77 51.18 86.46 94.60
Table 5: Ablation study on AdaMatch w/o AdaPatch.
CXR-to-Report Report-to-CXR
AdaPatch R@1 R@5 R@10 R@1 R@5 R@10
48.77 83.89 92.94 48.72 83.95 92.90
51.47 86.19 94.77 51.18 86.46 94.60
4.3 Ablation Study
Effectiveness of AdaPatch. To prove the effectiveness of AdaPatch, we compare the CXR-report retrieval performance of AdaMatch model with or without AdaPatch. In the AdaMatch model without AdaPatch, we use PVT-medium [43] as the image encoder. In Table 7, the AdaPatch module remarkably improves the R@1 by 2.70% and 2.46% for CXR-to-report and report-to-CXR retrieval, respectively, indicating the effectiveness of our AdaPatch.
Image Encoder with Different Stages. In AdaMatch, the image encoder consists of four stages, and the stage 2,3 and 4 include an AdaPatch module, respectively. To investigate the effectiveness of each stage for CXR-report alignment, we use patch embeddings from different stages to compute the CXR-report similarity for CXR-report retrieval task. In Table 6, the image encoder with three stages achieves the best retrieval performance with the R@1 of 51.47%, implying that the features of the 3rd stage are more beneficial for patch-word matching.
Table 6: CXR-report retrieval performance of AdaMatch with different stages.
CXR-to-Report Report-to-CXR
Stage R@1 R@5 R@10 R@1 R@5 R@10
2 42.28 78.75 89.85 42.81 78.96 89.67
3 51.47 86.19 94.77 51.18 86.46 94.60
4 49.68 83.04 92.67 49.39 83.09 92.84
Visualization of Adaptive Patches. To visualize adaptive patches in AdaPatch, we display several examples from the MIMIC-CXR datasets. In Fig. 8, each example includes the text, its text tokens, and the corresponding CXR image with the marked adaptive patches. The adaptive patches are marked by the bounding boxes with different colors. Each bounding box includes an arrow, which shows the shift of center from the fixed patch to the adaptive patch. In the first example, adaptive patches cover the pacemaker and its wire. Meanwhile, adaptive patches of the second example show the correct position of pleural effusions in the CXR image. These indicate that AdaPatch can correctly localize the region relevant to input text tokens.
Figure 8: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their shift of center from fixed patches.
5 Conclusion
We propose an AdaMatch for fine-grained image-text alignment, which presents the first work to adaptively associate image patches with words and improve the explainability of cyclic CXR-report generation. It includes the AdaPatch to acquire adaptive patches for abnormal regions, and performs patch-word alignment between adaptive patches with textual tokens. We implement cyclic CXR-report generation by using AdaMatch to provide explanations for the generation process. In addition, the fine-grained cyclic generation process supplies a natural explainability for the alignment between medical images and reports. Extensive experiments on two publicly available CXR datasets prove the effectiveness of our method and its superiority over previous methods.
Supplementary Material
In this supplementary material, we provide additional details to augment the content of the main paper. In Section 6, we conduct the ablation studies on the proposed AdaMatch-Cyclic. In Section 7, we demonstrate the visual results for CXR-to-report and report-to-CXR generation, and also visualize the adaptive patches and texts. In Section 8, we show implementation details of the proposed method.
6 Ablation Studies on AdaMatch-Cyclic
In the proposed AdaMatch-Cyclic, we leverage the AdaMatch to obtain keywords for CXR-to-report generation and keypatches for the report-to-CXR generation. To analyze the effectiveness of keywords and keypatches for the generation process, we ablate them in the AdaMatch-Cyclic on the MIMIC-CXR dataset to evaluate the CXR-report generation performance.
Effectiveness of keywords. As listed in Table 7, we ablate both keywords and keypatches of AdaMatch-Cyclic to create a baseline model. When we employ the keywords in the baseline model, the CXR-to-report generation performance improves significantly by about 0.03 in BLEU-4, indicating the effectiveness of the keywords obtained from AdaMatch. To further analyze the influence of the number of keywords $N_{w}$, we train the proposed AdaMatch-Cyclic with different $N_{w}=\left\{10,15,20\right\}$. In Table 8, AdaMatch-Cyclic achieves the best report generation performance with the ROUGE-L of 0.2859, when the $N_{w}$ is set to 10.
Effectiveness of keypatches. In Table 7, when we further apply keypatches to the baseline model with keywords, the CXR generation performance boosts remarkably by about 0.5 in FID score, implying the effectiveness of keypatches provided by AdaMatch. To investigate the influence of the number of keypatches $N_{p}$, we analyze the CXR generation performance when the number of keypatches ranges from 5 to 15. As shown in Table 9, AdaMatch-Cyclic achieves the best CXR generation performance with the FID score of 1.0916, when the number of keypatches is 5.
7 Visual Results
Table 7: The effectiveness of keywords and keypatches.
Keywords Keypatches BLEU-4 METEOR ROUGE-L FID
0.0778 0.1295 0.2402 1.5378
0.1059 0.1649 0.2813 1.5132
0.1060 0.1625 0.2859 1.0916
Table 8: The analysis on the different number of keywords ($N_{w}$).
$N_{w}$ BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L
10 0.3793 0.2346 0.1540 0.1060 0.1625 0.2859
15 0.3849 0.2265 0.1424 0.0937 0.1502 0.2668
20 0.3769 0.2281 0.1478 0.1001 0.1519 0.2762
Table 9: The analysis on the different number of keypatches($N_{p}$).
$N_{p}$ 5 10 15
FID 1.0916 1.6544 1.6720
Figure 9: Qualitative comparison with existing methods in CXR-to-report generation on the MIMIC-CXR (1st row) and OpenI (2nd row) datasets. The texts in different colors show similar meanings. The keywords on the right are obtained from AdaMatch model.
Figure 10: Visualization of the real and synthetic CXR images of the MIMIC-CXR (1st row) and the OpenI (2nd row) datasets.
Figure 11: Visualization of texts and the corresponding adaptive patches. The boxes and arrows in different colors show adaptive patches and their shift of center from fixed patches.
To prove the effectiveness of AdaMatch-Cyclic for CXR-to-report and report-to-CXR generation tasks, we demonstrate some visual results for both tasks. Moreover, to further visualize the explanation provided by AdaMatch, i.e. the adaptive patches and the corresponding texts.
CXR-to-Report Generation. As shown in Fig. 9, we compare CXR-to-report generation performance with existing methods on the MIMIC-CXR and OpenI datasets. In the first row, the proposed AdaMatch-Cyclic can capture the ‘Feeding tube tip’ and ‘mild bibasilar atelectasis’, while M2KT [47] cannot observe such device and lung abnormality. The second case of the OpenI dataset shows that ‘cardiac silhouette is mildly enlarged’ and ‘degenerative changes of the thoracic spine’ can be discovered by our method. These imply that our AdaMatch-Cyclic can generate a more comprehensive and complete medical report with the guidance of generated keywords in comparison to current methods.
Report-to-CXR Generation. Fig. 10 visualizes the real and synthetic CXR images of the MIMIC-CXR and OpenI datasets in comparison with existing methods. As depicted in the first example, the CXR image generated by AdaMatch-Cyclic shows ‘patchy opacity in the right lung’, while the CXR images generated by other methods do not include this pattern. In the second example, our AdaMatch-Cyclic can generate the CXR image with ‘low lung volumes’. These indicate the superiority of our AdaMatch-Cyclic over existing methods in report-to-CXR generation.
Adaptive Patches and Texts. In Fig. 11, we visualize adaptive patches in AdaPatch, the textual words, and the textual tokens for different cases from the MIMIC-CXR datasets. We highlight the adaptive patches with bounding boxes in different colors. Each bounding box has an arrow inside to show the shift of center from the fixed patch to the adaptive patch. In the first example, adaptive patches cover the pulmonary edema. Meanwhile, adaptive patches of the second example show the correct position of PICC device in the left lung of the CXR image. These suggest that AdaMatch-Cyclic can show the correspondence between the adaptive patches and textual words to provide the correct explanation for the CXR-report generation.
8 Implementation Details
In the AdaMatch-Cyclic model, we first train the AdaMatch model and then use the frozen AdaMatch to train LLM. The AdaMatch model consists of an image encoder and a text encoder. We utilize the PVT-medium [43] as the image encoder that includes AdaPatch module in stage 2, 3, and 4. The image encoder is pre-trained on the MIMIC-CXR dataset with the disease classification task. We adopt the pre-trained BioClinicalBERT [1] as text encoder. The image and text encoders are followed by two convolutional layers with batch normalization and the ReLU activation function, respectively, to reduce the feature dimension to 256. The patch size $s$ is set to 4 for stage 1, and 2 for stage 2, 3, and 4. The number of sampled feature points $m$ is 3. We optimize the AdaMatch using LAMB optimizer [51] ($\beta_{1}=0.9,\beta_{2}=0.999,\epsilon={10}^{-4}$). The cosine learning rate scheduler [29] with the base learning rate ($lr\_b$) of $6\times 10^{-3}$ is adopted to linearly warm up to the peak learning rate (i.e. $lr\_p=lr\_b\times\sqrt{\frac{bs\_all}{512}}$) during the first quarter of the total training epochs, where $bs\_all$ denotes the effective total batch size. The total training epoch is 15 and the per GPU batch size is 112. In AdaMatch-Cyclic, we use the pre-trained VQ-GAN [13] model to encode CXR images into image tokens and decode image tokens into CXR images. We adopt the dolly-v2-3b [9] model as the pre-trained LLM. The LLM has 5,1845 token types with the first 5,0821 token types for text tokens and the rest 1,024 token types for image tokens. We add 5,0821 to each image token value encoded by VQ-GAN. To decode the image tokens into images, we subtract 5,0821 from the image tokens generated by LLM and feed image tokens to VQ-GAN decoder to obtain the generated CXR images. We train the LLM with the AdamW [30] optimizer. The learning rate is initialized as $5\times{10}^{-6}$ and the total training epoch is 5. The per GPU batch size is set to 24. The hyper-parameters $\kappa_{0},\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4}$ are set as 200, 1, 1000, 20, and 5, respectively. In CXR-to-report generation, we extract keywords for each adaptive patch and use the keywords with the top 10 patch-word similarities as hints of LLM. In addition, we use 5 keypatches to guide the LLM in report-to-CXR generation. All the experiments are conducted on 8 Nvidia A100 40GB GPUs. The source code will be released.